OpenStudy (anonymous):

A plane flying with a tail wind flew at a speed of 460 mph, relative to the ground. When flying against the tailwind, it flew at a speed of 310 mph. Find the speed of the tailwind.

OpenStudy (anonymous):

(460-310)/2=75 so the answer is 75

OpenStudy (anonymous):

Thanks a bunch but can you explain it to me with the formula

OpenStudy (anonymous):

(460-310)/2 need to subtract 460-310 first then you divided by 2 then you will get 75

OpenStudy (anonymous):

Can I show you what I did to work it and maybe you can tell me where I went wrong? In class we have to use a formula like d = r(t)...x = flying with the tailwind, y = flying against the tailwind...or something like that...I am confused. I get 460 = x = y and under that I get 310 = x - y...but when I solve it all I get - 75 so I knew something was wrong...I just want to know what formula you used to plug in the 460 - 310/2....are you using formula r = d/t????

OpenStudy (anonymous):

Can anyone assist? I just need someone to explain why the equation to solve this word problem is r = d/t and why is 460 - 310 and why is it divided by 2? How is 2 substitute the t for time? I know we are trying to find out the rate...the tailwind but I need to grasp who and why it is solved this way or I won't be able to answer another problem similar to it.

OpenStudy (anonymous):

sorry how not who...

OpenStudy (anonymous):

I appreciate anyone who can explain this to me more...thank you in advance

OpenStudy (anonymous):

V tail = V + V wind V head = V - V wind -> V tail - V head = 2 V wind => V wind = 2 V wind/2

OpenStudy (anonymous):

Just think in common sense!

OpenStudy (anonymous):

Thank you Chlorophyll! That was helpful and exactly what I needed to understand the problem and how to solve it.

OpenStudy (anonymous):

My honor :)

OpenStudy (anonymous):

Understand that the velocity of plane different due to velocity of the wind!

OpenStudy (anonymous):

Yes I totally understand now since you broke it down like that. I could not even wrap my head around what the problem meant let alone how to solve it. Now I will know how to handle another word problem like it in the future.

OpenStudy (anonymous):

I wish the other askers would pursue their concern like you :)

OpenStudy (anonymous):

:) I do not like a quick answer. I like when someone takes the time to make sure I understand what I am solving. Math problems cannot be memorized...Math is the only subject I know that you either get it or you don't get it. I did not get it before but now I do! Again, I appreciate you taking the time to help me out tonight. I cannot move on with my school work when something is puzzling me. Take care!

OpenStudy (anonymous):

Hey, I'm the one pursue with the answer, it ends up that the askers running away, as usual :P